OpenStudy (anonymous):

Find the four smallest positive numbers theta such that tan(theta)=-1

OpenStudy (asnaseer):

in degrees these would be: 135, 315, 495, 675

OpenStudy (anonymous):

what about radians?

myininaya (myininaya):

multiply each of them by pi/(180 degrees)

OpenStudy (anonymous):

i dont understand

myininaya (myininaya):

\[135^o \cdot \frac{\pi}{180^o}\]

OpenStudy (asnaseer):

180 degrees = pi radians 1 degree = pi / 180 radians

myininaya (myininaya):

\[\frac{3 \pi}{4}\]

myininaya (myininaya):

for the first one

myininaya (myininaya):

all i did was multiply it by pi/(180 degrees)

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

\[\frac{3\pi}{4}\] and since tangent is periodic with period \[\pi\] you get the others by adding \[\pi\] repeatedly to \[\frac{3\pi}{4}\] and to hell with degrees.

myininaya (myininaya):

and i know you know what multiply means lol i just couldn't figure out what you couldn't understand about what i said

OpenStudy (anonymous):

couldn't figure out what degrees had to do with this problem

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

spend more time converting back and forth for no reason. the other day i saw someone write \[\sin(\frac{\pi}{4}+180)\]for petes sake!

myininaya (myininaya):

lol